I have this equation. I'm not sure if I did it right. This is what I have so far.
This is the equation:
0.5x+0.7=0.35x+1.3
0.85x+0.7+1.3
0.85+2

Answer by saifoo.khan(162)

Hello,
0.5x+0.7 = 0.35x + 1.3
0.5x - 0.35x = 1.3 - 0.7
0.15x = 0.6
x = 0.6/0.15
x = 60/15
x = 4
Answer..
Hope u got the answer.
Saifoo.khan
einstein.saifoo@gmail.com